Can someone answer number one? Thanks.
Answer #1

Data Given

\Delta T = 23^{0}C- 34^{0} C = -11^{0} C

Area A == 1.5 m2

Part A) h = 6 W/m2.0C, v = 0

\frac{\Delta Q}{\Delta t}= hA\Delta T = 6 W/m^{2}.^{0}C\times 1.5 m^{2}\times (-11^{0}) C= -99 W

Part B)

\frac{\Delta Q}{\Delta t}= hA\Delta T = 12 W/m^{2}.^{0}C\times 1.5 m^{2}\times (-11^{0}) C= -198 W

Part C)

\frac{\Delta Q}{\Delta t}= hA\Delta T = 28 W/m^{2}.^{0}C\times 1.5 m^{2}\times (-11^{0}) C= -462 W